OpenStudy (anonymous):

This is massively important, I've been stuck on the same lesson for almost a week now. A machine is set up to cut metal strips of varying lengths and widths based on the time (t) in minutes. The change in length is given by the function l(t)=t^2-√t and the change in width is given by w(t)=t^2+2t^1/2. Which function gives the change in area of the metal strips? A. a(t)=t^4+2t B. a(t)=t^4+2t+3t^5/2 C. a(t)=t^4−3t^5/2+2t D. a(t)=t^4+2t−2t^1/2+√t

OpenStudy (dumbcow):

Area = length*width \[(t^2 -\sqrt{t})*(t^2 +2\sqrt{t}) = t^4+2t^2 \sqrt{t}-t^2 \sqrt{t} -2t\] combine like terms \[t^4 +t^2 \sqrt{t} -2t\]

OpenStudy (anonymous):

Okay, but that's not one of the answers I have? Like, I'm not sure if I'm meant to be multiplying, or what? I'm sorry if this is a stupid question, I'm just totally lost.

OpenStudy (dumbcow):

yes area implies multiplying i dont know why it doesn't match, did i write the correct length and width functions?

OpenStudy (anonymous):

I think so? The answers I added are the multiple choice answers I have to choose from, so which one looks like the answer?

OpenStudy (dumbcow):

the one i wrote looks like the answer lol http://www.wolframalpha.com/input/?i=%28t%5E2+-sqrt%28t%29%29%28t%5E2+%2B2+sqrt%28t%29%29

OpenStudy (anonymous):

So which letter of the multiple choice then? It looks like the answer to me too, but it's got to be one of the multiple choice, and maybe I'm just missing something in the way you formatted your answer.